George Madrid 14 Saint Paul Street, #2 Cambridge, MA 02139

(617) 354-7853

June 19, 1995

Dear Members of the Committee on Academic Performance,

I would like to address the circumstances regarding my performance during the Spring term of 1995. Personal issues of which the CSS is aware were detrimental to both my academic and professional life. I spoke to my professors and took action which I considered appropriate and sufficient to salvage the term. I was surprised and confused when I received the letter from the CAP and more surprised when I received my grade report. I have discussed my situation with Dean Henderson of Counseling and Support Services, and he is aware of the issues involved.

In addition to being a student at MIT, I am a full-time employee at a local computer company called Dimensional Insight, Inc. I work to pay my tuition. I get no aid from my parents. I enrolled last term as a light-load student as I had for Fall 1994. I was not concerned about the work load since I had successfully passed all my classes in the Fall with no trouble. In February, I was enthusiastic and looking forward to an enjoyable and successful term.

I have been dating a young woman since late October. She had an extremely difficult term this Spring. The Dean’s Office is aware of her problems which has an extreme impact on her school life. Every time she just about caught up, another difficulty arose, unrelated to the previous and through no fault of her own.

I mention her difficulties because they are linked to mine. In being supportive to her, my own work suffered. I tried not to let her realize how much of an effect she was having on me, since she needed someone and resisted my attempts to send her to CSS or to the psychiatrists in the MIT Medical Department. I fell behind both at work and at school.

At this point, I approached both my professors to discuss my absences with them. As a result, I dropped my French class. I remained in my math class. After my discussion with Professor Sacks, I believed that if I turned in the homework for the classes which I missed, I would be okay. There were no exams or finals in this class, so that seemed more than reasonable. Since I was still facing a great deal of pressure at work, I wasn’t able to do this immediately, but when I met my work deadlines, I did the homework. This was still before the end of the term.

Since I thought I had done what Professor Sacks asked me, I was very surprised by the events of last week. I received the letter from the CAP last Wednesday and my grade report on Thursday. Since that time, I have tried to reach Professor Sacks but I have failed. He has not responded to my email and has not been in his office when I have looked for him. I believe that the grade is just a misunderstanding of some sort. Perhaps my name and situation slipped his mind. Perhaps I wasn’t clear on his instructions. I don’t know.

My current plans are to quit my job in August and return to MIT full-time until I have completed my degree. I am looking forward to being a student unencumbered by the responsibilities or a full-time job. I am confident that without the additional pressures of a full-time job, I can finish up in a timely manner and proceed with the rest of my life.

With regards,

George Madrid